Example: Types of Float . You have $3,000 in your checking account. You just deposited $2,000 and wrote a check for $2,500. - What is the disbursement float? - What is the collection float? - What is the net float? - What is your book balance? What is your available balance? all rignts
Answer #1

For the check written:

Disbursement float = Available Balance - book Balance

Disbursement float = 3,000 - (3,000 - 2,500)

Disbursement float = 2,500

Part 2 - Collection Float

For the check deposited:

Collection Float = Available Balance - book Balance

Collection Float = 3,000 - (3,000 + 2,000)

Collection Float = -2,000

Part 3 - Net Float

Net Float = Disbursement float + Collection Float

Net Float = 2,500 - 2,000

Net Float = 500

Part 4:

Book Balance = Beginning Book balance + Cheque recieved - Cheque issued

Book Balance = 3,000 + 2,000 - 2,500

Book Balance = 2,500

Part 5

Available balance = 3,000